OpenStudy (anonymous):

Please help important test question algebra 1 1. If -nx^2 + tx + c = 0, what is x equal to?

OpenStudy (anonymous):

Isolate N.

OpenStudy (anonymous):

2. The graphs of f(x) and g(x) are shown below: If f(x) = (x + 7)2, which of the following is g(x) based on the translation? A. g(x) = x2 + 18x + 81 b. g(x) = x2 + 10x + 25 C. g(x) = x2 + 14x + 47 D. g(x) = x2 + 14x + 51

OpenStudy (anonymous):

@thebossman how?

OpenStudy (anonymous):

I got answer one just need help with question 2 someone please help

OpenStudy (anonymous):

i don't understand what the question is asking

OpenStudy (anonymous):

you have two graphs there one looks like the graph of \(y=(x+7)^2\) and the other is the graph of \(y=(x+9)^2\)

OpenStudy (zzr0ck3r):

I dont understand which question is the one being asked

OpenStudy (anonymous):

ooh i see !!

OpenStudy (anonymous):

your job it to multiply out \[g(x)=(x+9)^2\] and see which of the above 4 answers you get shouldn't take more than five seconds a guess will do

OpenStudy (anonymous):

Im confused how do i multiply that out? what do i put in for x

OpenStudy (anonymous):

ok i got g(x) = x2 + 18x + 81TY